What is 1 inch minus 1/2 inches?
1
Expert's answer
2013-03-12T13:24:55-0400
1 inch minus 1/2 inches = (1 - 1/2 ) inches
1 - 1/2 = | 1 = 2/2 | =2/2 - 1/2 = (2-1)/2 = 1/2
Therefore, 1 inch minus 1/2 inches= (1 - 1/2 ) inches = 1/2 inches
Answer: 1 inchminus 1/2 inches = 1/2 inches